OpenStudy (xxorejina):

The car at each vertex of a Ferris wheel holds a maximum of 5 people. The sum of the measures of the angles of the Ferris wheel is 7740º. What is the maximum number of people that the Ferris wheel can hold? 250 225 300 200

OpenStudy (aihberkhan):

Okay. The sum of interior angles of a polygon = (n-2)*180°, where n is number of sides / vertices We already have this sum, so we will go backwards to find n. 7740 = 180(n-2) Divide by 180 43 = n - 2 45 = n So we have 45 vertices. Each vertex can hold 5 people, hence, 45 * 5 = 225 people all together.

OpenStudy (aihberkhan):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @xxorejina

OpenStudy (anonymous):

me to